# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ E.3 Design Principles for Spectral Weighting
Spectral weighting controls the direction-specific amplification of Anti-regularization. The key goal is to suppress over-amplification along directions with small singular or eigenvalues.
Principle 1: Maintain safety margin
Always enforce the generalized safety condition $\lambda<1/\sigma_{\max}(A)$. In practice, apply a margin:
@@
Under simultaneous diagonalization, this becomes $\lambda\leq(1-\epsilon)\min_{j}\sigma_{j}/w_{j}$.
Principle 2: Monotonic increasing weights
To avoid excessive amplification in small-$\sigma_{j}$ directions, set $w_{j}$ as a non-decreasing function of $\sigma_{j}$. Typical designs include:
@@
If $W=c\,\hat{\Sigma}$, then $A=cI$, yielding $\operatorname{tr}(S_{\lambda}^{(W)})=\tfrac{r}{1-\lambda c}$, i.e., uniform amplification across all directions with minimal spectral distortion.
Principle 3: Degrees-of-freedom budgeting
Target a desired per-sample complexity $\kappa>0$ via
$\frac{1}{n}\operatorname{tr}(S_{\lambda}^{(W)})=\kappa,\quad\text{or equivalently}\quad\sum_{j}\frac{1}{1-\lambda a_{j}}=\kappa n.$
Solve numerically, clipping $\lambda\leq(1-\epsilon)/\sigma_{\max}(A)$ to prevent excessive sensitivity in small-$a_{j}$ directions.
Principle 4: Suppress noisy directions
To suppress amplification in directions known to be noisy or unstable, reduce $w_{j}$ accordingly. For example, use leverage scores $\ell_{j}$ or cross-validation-based noise estimates $\widehat{\mathrm{Var}}_{j}$:
$w_{j}\propto\frac{\sigma_{j}}{\sigma_{j}+\tau_{\text{dir}}^{2}\,\widehat{\mathrm{Var}}_{j}}$
so that larger noise implies smaller $w_{j}$, thereby suppressing $\tfrac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}}$ expansion.
Principle 5: Hybrid schedule
@@
$\lambda(n)=\lambda(n_{0})\left(\frac{n_{0}}{n}\right)^{\alpha},\qquad\alpha\geq 1,$
while keeping $W$ fixed or slowly adapting (e.g., via moving averages over $w_{j}$). This stabilizes $\operatorname{tr}(S_{\lambda}^{(W)})/n$ over time.
Checklist
  • (i)
    Estimate $\sigma_{\max}(A)$, then clip $\lambda$ accordingly.
  • (ii)
    Monitor $\operatorname{tr}(S_{\lambda}^{(W)})$ and $\operatorname{tr}((S_{\lambda}^{(W)})^{2})$.
  • (iii)
    Reduce $w_{j}$ for small $\sigma_{j}$ directions to mitigate amplification.
  • (iv)
    Consider using $W=c\hat{\Sigma}$ as an approximation to improve $A$’s condition number.
Appendix F Classification: Lack of Lower Bound and Remedies
@@ F.1 Constructive Examples of Unboundedness
We demonstrate that on separable data, the cross-entropy $H$ with negative $L_{2}$ reward
@@
can become unbounded from below.
Proposition 25 (Binary Logistic: No Lower Bound).
Let $y_{i}\in\{\pm 1\}$, logit $z_{i}(\theta)=\theta^{\top}x_{i}$, and loss $\ell_{i}(\theta)=\log(1+\exp(-y_{i}z_{i}(\theta)))$. If the data is linearly separable, i.e., there exists $u\neq 0$ and $\delta>0$ such that $y_{i}\,u^{\top}x_{i}\geq\delta$ for all $i$, then for the trajectory $\theta_{t}=tu$,
@@
Each term satisfies $\ell_{i}(\theta_{t})\leq\exp(-t\delta)\to 0$, while the reward term diverges as $\frac{\lambda}{2}\|\theta_{t}\|^{2}=-\frac{\lambda}{2}t^{2}\|u\|^{2}\to-\infty$. Thus, $\hat{F}_{\lambda}(\theta_{t})\to$-$\infty$. ∎
Proposition 26 (Multiclass Softmax: No Lower Bound).
For multiclass logits $f_{\theta}(x)\in\mathbb{R}^{C}$, define the correct-class margin $m_{\theta}(x_{i},y_{i})=f_{\theta}(x_{i})_{y_{i}}-\max_{k\neq y_{i}}f_{\theta}(x_{i})_{k}$. If there exists $u$ and $\delta>0$ such that $m_{tu}(x_{i},y_{i})\geq t\delta$ for all $i$, then
$\frac{1}{n}\sum_{i}H\!\big{(}p_{y_{i}},q_{tu}(x_{i})\big{)}\to 0,\quad\text{but}\quad-\frac{\lambda}{2}\|tu\|^{2}\to-\infty,$
and hence $\hat{F}_{\lambda}(tu)\to$-$\infty$.
Remark 11 (Interpretation).
Classification loss $H$ can approach zero on separable data by scaling $\|\theta\|$ large, but the negative $L_{2}$ reward $\tfrac{\lambda}{2}\|\theta\|^{2}$ diverges quadratically. Thus, negative $L_{2}$ regularization alone is insufficient to ensure stability in classification.
F.2 Trust Region Remedy
@@
Let $\Theta_{B}=\{\theta:\ \|\theta\|\leq B\}$ be a trust region of radius $B>0$. Consider the constrained problem
@@
If $H(\cdot)$ is continuous, then $\hat{F}_{\lambda}$ is continuous on $\Theta_{B}$, and by the Weierstrass theorem, a minimizer $\hat{\theta}_{\lambda,B}$ exists.
@@
$\Theta_{B}$ is compact (closed and bounded subset of $\mathbb{R}^{p}$), and $\hat{F}_{\lambda}$ is continuous. The Weierstrass theorem guarantees the existence of a minimum. ∎
Remark 12 (Practical Guidelines).
The radius $B$ may be chosen based on validation performance or numerical stability (e.g., logit saturation, gradient explosion). In practice, this is implemented via parameter clipping (e.g., $\|\theta\|>B\Rightarrow\theta\leftarrow B\,\theta/\|\theta\|$), or by weight normalization or norm constraints.
F.3 Bounded Margin Reward $\phi$
Instead of negative $L_{2}$, we use a bounded margin reward $\phi$ to guarantee a lower bound.
@@ F.3 Bounded Margin Reward $\phi$
Hence, $\hat{F}_{\lambda}^{(\phi)}$ is bounded below.
@@
Since $H\geq 0$ and $\phi\leq\phi_{\max}$, $\hat{F}_{\lambda}^{(\phi)}(\theta)\geq-\lambda\cdot\tfrac{1}{n}\sum_{i}\phi_{\max}=-\lambda\phi_{\max}$. ∎
Remark 13 (Existence of Minimizer).
A lower bound does not guarantee the existence of a minimizer (e.g., when limit values are reached only at infinity). Thus, adding a trust region (see §F.2) or clipping logits/parameters (e.g., $|f_{\theta}(x)_{k}|\leq M$) to compactify level sets is advised. In practice, combining a bounded $\phi$ with parameter or logit clipping is recommended.
Design Examples
The following $\phi$ functions satisfy boundedness and Lipschitz conditions:
@@
If the network is Lipschitz in parameters with constant $G$ (i.e., $|m_{\theta}(x,y)-m_{\theta^{\prime}}(x,y)|\leq G\|\theta-\theta^{\prime}\|$), and $\phi$ is $L_{\phi}$-Lipschitz, then
@@
so the gradient norm of the reward term is bounded by $\lambda L_{\phi}G$, suppressing numerical explosions. Combining with logit/gradient clipping further enhances stability.
@@
(i) Calibrate $\phi_{\max},\gamma$ on validation set (prevent over-rewarding large margins);
(ii) Combine logit clipping $f_{\theta}(x)\in[-M,M]^{C}$ and parameter clipping $\|\theta\|\leq B$;
(iii) Use sample-dependent scaling $\lambda(n)=\lambda(n_{0})(n_{0}/n)^{\alpha}$, $\alpha\geq\tfrac{1}{2}$;
(iv) For non-separable data, use large $\gamma$ to avoid early saturation of $\phi$.
F.4 Generalization Bound Derivation
We present a generalization upper bound on 0–1 risk when using a bounded margin reward $\phi$ in classification. We adopt the same notation as in §F.3 and assume the following:
Assumptions
(i) Either a trust region $\Theta_{B}=\{\theta:\|\theta\|\leq B\}$ or logit clipping $|f_{\theta}(x)_{k}|\leq M$ ensures compactness at the parameter/logit level.
(ii) The network is $G$-Lipschitz in parameters: $|m_{\theta}(x,y)-m_{\theta^{\prime}}(x,y)|\leq G\|\theta-\theta^{\prime}\|$.
(iii) The reward $\phi:\mathbb{R}\to[0,\phi_{\max}]$ is $L_{\phi}$-Lipschitz.
(iv) A standard margin condition with margin width $\gamma>0$ holds via ramp/hinge surrogate losses.
Theorem 29 (Combined Margin–Lipschitz Generalization Bound).
Under the above assumptions, for any $\delta\in(0,1)$, with probability at least $1-\delta$, the following holds for all $\theta\in\Theta_{B}$:
$\mathcal{R}_{0\text{-}1}(f_{\theta})\;\leq\;\frac{1}{n}\sum_{i=1}^{n}H\!\big{(}p_{y_{i}},q_{\theta}(x_{i})\big{)}\;-\;\lambda\,\frac{1}{n}\sum_{i=1}^{n}\phi\!\big{(}m_{\theta}(x_{i},y_{i})\big{)}\;+\;C_{1}\,\frac{G}{\gamma}\sqrt{\frac{1}{n}}\;+\;C_{2}\,|\lambda|\,L_{\phi}\,\sqrt{\frac{1}{n}}\;+\;C_{3}\,\sqrt{\frac{\log(1/\delta)}{n}}.$
Here, $C_{1},C_{2},C_{3}>0$ are universal constants that may depend on the data radius or clipping constants.
@@ F.4 Generalization Bound Derivation
We outline the proof using the following steps:
@@
    Reduction to Margin.
    $\mathcal{R}_{0\text{-}1}(f_{\theta})=\mathbb{E}\big{[}\mathbf{1}\{m_{\theta}(x,y)\leq 0\}\big{]}\ \leq\ \mathbb{E}\big{[}L_{\gamma}(m_{\theta}(x,y))\big{]},$
    where $L_{\gamma}$ is a ramp/hinge-type surrogate satisfying $\mathbf{1}\{m\leq 0\}\leq L_{\gamma}(m)\leq\mathbf{1}\{m\leq\gamma\}$.
@@
    Empirical–Expected Deviation (Rademacher/Contraction).
    Since $L_{\gamma}\circ m_{\theta}$ is $(G/\gamma)$-Lipschitz:
    $\mathbb{E}\big{[}L_{\gamma}(m_{\theta})\big{]}\ \leq\ \frac{1}{n}\sum_{i=1}^{n}L_{\gamma}(m_{\theta}(x_{i},y_{i}))+C_{1}\,\frac{G}{\gamma}\sqrt{\frac{1}{n}}.$
@@
    Surrogate–CE Loss Comparison.
    Via standard calibration arguments, for some constant $c_{0}\geq 0$:
@@
    Complexity of Reward Term.
    Since $\phi\circ m_{\theta}$ is $(L_{\phi}G)$-Lipschitz:
    $\Bigg{|}\frac{1}{n}\sum_{i=1}^{n}\phi\!\big{(}m_{\theta}(x_{i},y_{i})\big{)}-\mathbb{E}\big{[}\phi(m_{\theta}(x,y))\big{]}\Bigg{|}\ \leq\ C_{2}\,L_{\phi}\,G\,\sqrt{\frac{1}{n}}.$
    Hence, the $\lambda\,\tfrac{1}{n}\sum\phi$ term contributes a correction of $C_{2}\,|\lambda|\,L_{\phi}\,\sqrt{1/n}$ to the bound.
@@
    Probability Boosting and Final Bound.
    Using McDiarmid’s or Bernstein’s inequality, we obtain with probability $1-\delta$ an additional deviation term $C_{3}\,\sqrt{\log(1/\delta)/n}$, yielding
    $\mathcal{R}_{0\text{-}1}(f_{\theta})\ \leq\ \frac{1}{n}\sum_{i=1}^{n}H(p_{y_{i}},q_{\theta}(x_{i}))-\lambda\,\frac{1}{n}\sum_{i=1}^{n}\phi(m_{\theta}(x_{i},y_{i}))+C_{1}\,\frac{G}{\gamma}\sqrt{\frac{1}{n}}+C_{2}\,|\lambda|\,L_{\phi}\,\sqrt{\frac{1}{n}}+C_{3}\,\sqrt{\frac{\log(1/\delta)}{n}}.$
Constants $C_{1},C_{2},C_{3}$ are universal and may depend on the input norm bound or clipping radius. ∎
@@
The data-dependent terms in the bound consist of the empirical loss term $\tfrac{1}{n}\sum H-\lambda\tfrac{1}{n}\sum\phi$, while the complexity penalties separate into $O\!\big{(}\tfrac{G}{\gamma\sqrt{n}}\big{)}$ and $O\!\big{(}|\lambda|L_{\phi}/\sqrt{n}\big{)}$. Thus, to preserve generalization balance, $|\lambda|$ should decay at least as fast as $n^{-1/2}$.
F.5 Proof of Sample-Size Scaling Rule
Theorem 30 (Sample-Size Scaling Rule in Classification).
Let $\lambda(n)=\lambda(n_{0})\left(\tfrac{n_{0}}{n}\right)^{\alpha}$. If $\alpha\geq\tfrac{1}{2}$, then the complexity penalty in Theorem 29 remains bounded by $O(n^{-1/2})$. In particular, when $\alpha>\tfrac{1}{2}$, the reward-induced complexity term $C_{2}|\lambda|L_{\phi}/\sqrt{n}$ decays faster, i.e., $o(n^{-1/2})$.

### v2

@@ E.3 Design Principles for Spectral Weighting
Spectral weighting regulates the directional amplification of Anti-regularization. The key objective is to suppress excessive amplification in directions associated with small singular values or eigenvalues.
Principle 1: Maintain a safety margin.
Always satisfy the generalized safety condition $\lambda<1/\sigma_{\max}(A)$. In practice, impose
@@
Under simultaneous diagonalization, this reduces to $\lambda\leq(1-\epsilon)\min_{j}\sigma_{j}/w_{j}$.
Principle 2: Monotone increasing weights.
To prevent over-amplification in small-$\sigma_{j}$ directions, set $w_{j}$ as a non-decreasing function of $\sigma_{j}$. Representative choices include
@@
When $W=c\,\hat{\Sigma}$, then $A=cI$ and $\operatorname{tr}(S_{\lambda}^{(W)})=\tfrac{r}{1-\lambda c}$, so that all directions are uniformly amplified, minimizing spectral distortion.
Principle 3: Control the degrees-of-freedom budget.
For a target $\kappa>0$, enforce $\tfrac{1}{n}\operatorname{tr}(S_{\lambda}^{(W)})=\kappa$. With the whitened representation,
$\frac{1}{n}\sum_{j}\frac{1}{1-\lambda a_{j}}=\kappa,$
which can be solved numerically. As $\lambda$ increases, sensitivity to small $a_{j}$ grows sharply, so couple Newton updates with clipping at $(1-\epsilon)/\sigma_{\max}(A)$.
Principle 4: Suppress noisy directions.
If certain feature directions are noisy or unstable, reduce the corresponding $a_{j}$ by setting smaller $w_{j}$. For example, based on leverage scores $\ell_{j}$ or variance estimates $\widehat{\mathrm{Var}}_{j}$, set
$w_{j}\propto\frac{\sigma_{j}}{\,\sigma_{j}+\tau_{\text{dir}}^{2}\,\widehat{\mathrm{Var}}_{j}\,}$,
so that noisy directions receive smaller weights, thereby suppressing the inflation of $\tfrac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}}$.
Principle 5: Hybrid schedules.
@@
$\lambda(n)=\lambda(n_{0})\Big{(}\frac{n_{0}}{n}\Big{)}^{\alpha},\qquad\alpha\geq 1,$
while fixing $W$ or adapting it slowly (e.g., moving average of $w_{j}$) to keep $\operatorname{tr}\!\big{(}S_{\lambda}^{(W)}\big{)}/n$ stable.
Checklist.
[…]
Appendix F Classification: Absence of Lower Bound and Remedies
@@ F.1 Constructive examples of unboundedness
This section demonstrates that, for separable data, the combination of cross-entropy $H$ and a negative $L_{2}$ reward
@@
may lack a finite lower bound.
Proposition 26 (Binary logistic regression: absence of lower bound).
Let binary labels $y_{i}\in\{\pm 1\}$, logits $z_{i}(\theta)=\theta^{\top}x_{i}$, and losses $\ell_{i}(\theta)=\log(1+\exp(-y_{i}z_{i}(\theta)))$. If the data are linearly separable so that there exists $u\neq 0$ and $\delta>0$ with $y_{i}\,u^{\top}x_{i}\geq\delta\ \forall i$, then along the trajectory $\theta_{t}=t\,u$,
@@
For each term, $\ell_{i}(\theta_{t})=\log(1+\exp(-t\,y_{i}u^{\top}x_{i}))\leq\exp(-t\delta)$, hence $\tfrac{1}{n}\sum_{i}\ell_{i}(\theta_{t})\leq\exp(-t\delta)\to 0$. Meanwhile, the reward term is $\tfrac{\lambda}{2}\|\theta_{t}\|^{2}=-\tfrac{\lambda}{2}t^{2}\|u\|^{2}\to-\infty$. Thus, $\hat{F}_{\lambda}(\theta_{t})\to$-$\infty$. ∎
Proposition 27 (Multiclass softmax: absence of lower bound).
In the multiclass case, let the logit vector be $f_{\theta}(x)\in\mathbb{R}^{C}$, with margin $m_{\theta}(x_{i},y_{i})=f_{\theta}(x_{i})_{y_{i}}-\max_{k\neq y_{i}}f_{\theta}(x_{i})_{k}$. If there exists $u$ and $\delta>0$ such that for all $i$, $m_{tu}(x_{i},y_{i})\geq t\delta$, then
$\frac{1}{n}\sum_{i}H\!\big{(}p_{y_{i}},q_{tu}(x_{i})\big{)}\to 0,\quad\text{while}\quad-\frac{\lambda}{2}\|tu\|^{2}\to-\infty,$
so that $\hat{F}_{\lambda}(tu)\to$-$\infty$.
Remark 15 (Implication).
For separable data, the classification loss $H$ can be driven arbitrarily close to zero by scaling up $\|\theta\|$. However, the negative $L_{2}$ reward contributes a term $\tfrac{\lambda}{2}\|\theta\|^{2}$ that diverges quadratically, so the overall objective descends to -$\infty$. Thus, negative $L_{2}$ alone cannot guarantee stability in classification.
F.2 Trust-region remedy
@@
Consider a trust region $\Theta_{B}=\{\theta:\ \|\theta\|\leq B\}$ with radius $B>0$, and the optimization problem
@@
If $H(\cdot)$ is continuous, then $\hat{F}_{\lambda}$ is continuous on $\Theta_{B}$. By the Weierstrass theorem, a minimizer $\hat{\theta}_{\lambda,B}$ exists.
@@
The set $\Theta_{B}$ is compact in $\mathbb{R}^{p}$ (closed and bounded), and $\hat{F}_{\lambda}$ is continuous. By the Weierstrass theorem, the minimum is attained. ∎
Remark 16 (Selection guidelines).
The radius $B$ can be chosen based on validation performance or numerical stability criteria, such as avoiding logit saturation or gradient explosion. In practice, trust regions can be enforced via parameter clipping, i.e., rescaling $\theta\leftarrow B\,\theta/\|\theta\|$ whenever $\|\theta\|>B$, or through weight normalization and norm-constrained parameterizations.
F.3 Upper-bounded margin reward $\phi$
Instead of a negative $L_{2}$ penalty, one can employ an upper-bounded margin reward $\phi$ to guarantee the existence of a lower bound.
@@ F.3 Upper-bounded margin reward $\phi$
In particular, $\hat{F}_{\lambda}^{(\phi)}$ admits a finite lower bound.
@@
Since $H\geq 0$ and $\phi\leq\phi_{\max}$, $\hat{F}_{\lambda}^{(\phi)}(\theta)\geq 0-\lambda\cdot\tfrac{1}{n}\sum_{i}\phi_{\max}=-\lambda\phi_{\max}$. ∎
Remark 17 (Existence of minimizers).
A lower bound alone does not guarantee the existence of a minimizer, as the infimum may be attained only at infinity. To ensure existence, one may impose a trust region (§F.2), or apply parameter or logit clipping (e.g., $|f_{\theta}(x)_{k}|\leq M$) to make the level sets compact. In practice, it is recommended to combine bounded $\phi$ with parameter or logit clipping.
Example designs
The following choices of $\phi$ satisfy boundedness and Lipschitz conditions:
@@
If the network is $G$-Lipschitz with respect to the parameters (i.e., $|m_{\theta}(x,y)-m_{\theta^{\prime}}(x,y)|\leq G\|\theta-\theta^{\prime}\|$) and $\phi$ is $L_{\phi}$-Lipschitz, then
@@
so the gradient magnitude of the reward term is controlled by $\lambda L_{\phi}G$, mitigating numerical blow-up. Additional logit and gradient clipping further reinforce stability.
@@
(i) Calibrate $\phi_{\max},\gamma$ on a validation set (to avoid over-rewarding large margins);  (ii) combine with logit clipping $f_{\theta}(x)\in[-M,M]^{C}$ and parameter clipping $\|\theta\|\leq B$;  (iii) use a decay schedule $\lambda(n)=\lambda(n_{0})(n_{0}/n)^{\alpha}$ with $\alpha\geq\tfrac{1}{2}$;  (iv) for non-separable data, set the saturation threshold $\gamma$ larger to prevent premature saturation of the reward.
Appendix G Classification: Generalization bound and $|\lambda|$ scaling
G.1 Derivation of the generalization bound
This section derives a 0–1 risk bound for classification when an upper-bounded margin reward $\phi$ is used. The notation follows §F.3, with the following assumptions:
Assumptions.
(i) A trust region $\Theta_{B}=\{\theta:\|\theta\|\leq B\}$ or logit clipping $|f_{\theta}(x)_{k}|\leq M$ ensures compactness at the parameter/logit level. (ii) The network is $G$-Lipschitz with respect to the parameters: $|m_{\theta}(x,y)-m_{\theta^{\prime}}(x,y)|\leq G\|\theta-\theta^{\prime}\|$. (iii) The margin reward $\phi:\mathbb{R}\to[0,\phi_{\max}]$ is $L_{\phi}$-Lipschitz. (iv) A standard margin bound (width $\gamma>0$) via ramp/hinge transformations holds.
Theorem 30 (Margin–Lipschitz combined generalization bound).
Under the above assumptions, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $\theta\in\Theta_{B}$,
$\mathcal{R}_{0\text{-}1}(f_{\theta})\;\leq\;\frac{1}{n}\sum_{i=1}^{n}H\!\big{(}p_{y_{i}},q_{\theta}(x_{i})\big{)}-\lambda\,\frac{1}{n}\sum_{i=1}^{n}\phi\!\big{(}m_{\theta}(x_{i},y_{i})\big{)}+C_{1}\frac{G}{\gamma}\sqrt{\frac{1}{n}}+C_{2}\,|\lambda|\,L_{\phi}\,\sqrt{\frac{1}{n}}+C_{3}\sqrt{\frac{\log(1/\delta)}{n}},$
where $C_{1},C_{2},C_{3}>0$ are universal constants that may depend on data radius or clipping constants.
@@ G.1 Derivation of the generalization bound
We outline the main steps:
@@
    Reduction to a margin surrogate. Using a ramp/hinge surrogate $L_{\gamma}$,
    $\mathcal{R}_{0\text{-}1}(f_{\theta})=\mathbb{E}[\mathbf{1}\{m_{\theta}(x,y)\leq 0\}]\ \leq\ \mathbb{E}[L_{\gamma}(m_{\theta}(x,y))].$
@@
    Empirical–expectation gap (Rademacher contraction). Since $L_{\gamma}\circ m_{\theta}$ is $(G/\gamma)$-Lipschitz,
    $\mathbb{E}[L_{\gamma}(m_{\theta})]\ \leq\ \frac{1}{n}\sum_{i=1}^{n}L_{\gamma}(m_{\theta}(x_{i},y_{i}))+C_{1}\frac{G}{\gamma}\sqrt{\frac{1}{n}}.$
@@
    Comparison with cross-entropy. By calibration/upper-bound arguments, for some constant $c_{0}$,
@@
    Complexity of the reward term. Since $\phi\circ m_{\theta}$ is $(L_{\phi}G)$-Lipschitz,
    $\Bigg{|}\frac{1}{n}\sum_{i=1}^{n}\phi(m_{\theta}(x_{i},y_{i}))-\mathbb{E}[\phi(m_{\theta}(x,y))]\Bigg{|}\ \leq\ C_{2}\,L_{\phi}\,G\,\sqrt{\frac{1}{n}}.$
    Thus the empirical reward contributes an additional deviation term $C_{2}\,|\lambda|\,L_{\phi}\,\sqrt{1/n}$.
@@
    Probabilistic concentration. Applying McDiarmid’s or Bernstein’s inequality adds the final term $C_{3}\sqrt{\log(1/\delta)/n}$.
Combining all steps yields the stated inequality. ∎
@@
The data-dependent part of the bound is the empirical term $\tfrac{1}{n}\sum H-\lambda\tfrac{1}{n}\sum\phi$. The complexity penalties separate into $O(G/(\gamma\sqrt{n}))$ and $O(|\lambda|L_{\phi}/\sqrt{n})$. Therefore, to maintain balance, $|\lambda|$ must decay at least on the order of $n^{-1/2}$.
G.2 Proof of the sample-size scaling rule
Theorem 31 (Sample-size scaling rule for classification).
Let $\lambda(n)=\lambda(n_{0})\big{(}\tfrac{n_{0}}{n}\big{)}^{\alpha}$. If $\alpha\geq\tfrac{1}{2}$, then the complexity penalties in Theorem 30 remain $O(n^{-1/2})$. In particular, when $\alpha>\tfrac{1}{2}$, the reward-induced complexity term $C_{2}|\lambda|L_{\phi}/\sqrt{n}$ decays as $o(n^{-1/2})$.